OpenStudy (anonymous):

If Jim takes 6 days to paint a house alone and Mike takes 8 days to paint the same house alone, what part of the job will be completed if both boys work for 2 days?

OpenStudy (anonymous):

here, we compare the rates of work by individual person. rate of work done by Jim = \(\Large {1\over6}\) per day rate of work done by Mike = \(\Large {1\over8}\) per day together, their rate will be \(\Large{{1\over6}+{1\over8}=?}\) perday

OpenStudy (anonymous):

that will be the amount of work done in "1" day in "2" days, the amount of work done will be twice of that

OpenStudy (anonymous):

Ahh I see. Thanks very much.

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

jim takes 6 days to paint the house thus his 1 day work=1/6 mike take 8 days to complete the work thus his I day work=1/8 thus jim's and mike's 1 day work together=1/6+1/8 =7/28 total work is completed by jim and mike in days=1/7/28 =4 days let the total work be 1 work completed in 4 days= 1 work completed in 1 day= 1/4 work completed in 2 days=1/4*2 =1/2 Ans- Jim and Mike will complete half of the work if they will work together for 2 days.